OpenStudy (anonymous):

I NEED HELP.please does anybody know the answer to this? Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2)

OpenStudy (anonymous):

First you're gonna need to find the slope. Do you know how to do that?

OpenStudy (anonymous):

y2-y1/x2-x1 ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

i get -10/20

OpenStudy (anonymous):

Yep, which simplifies to -1/2 now, point slope form looks like this: y-y1=m(x-x1) M=slope (x1, y1)= one of the points you wanna use

OpenStudy (anonymous):

so is the answer (y-8)=-1/2 (x+17)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ahh thank you so much!!

OpenStudy (anonymous):

You're welcome